OpenStudy (erinweeks):

(1.08)Find x and m QRP if m QRP = 2x – 1, m PRS = 3x + 2 and m QRS = 86. Is the angle bisector of QRS? Support your answer.

OpenStudy (erinweeks):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

because angles QRP and PRS combine to form angle QRS, this means you can say (angle QRP) + (angle PRS) = angle QRS (2x-1) + (3x+2) = 86 ... make the proper substitutions now solve for x

jimthompson5910 (jim_thompson5910):

I'll let you do that and tell me what you get

OpenStudy (erinweeks):

2x-1+3x+2=86 2x+1+3x=86 5x+1=86 -1 -1 5x=85 x=17

OpenStudy (erinweeks):

This is what I put on my assignment and her feed back was You didn’t answer the question. Is ray RP the angle bisector of angle QRS?

jimthompson5910 (jim_thompson5910):

so if x = 17, then what are angles QRP and PRS?

OpenStudy (erinweeks):

QRP=2(17)-1=33 PRS=3(17)+2=53

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (erinweeks):

Then leaves the question Is ray RP the angle bisector of angle QRS?

jimthompson5910 (jim_thompson5910):

so, what does that tell us?

jimthompson5910 (jim_thompson5910):

what does it mean if something bisects an angle?

OpenStudy (erinweeks):

That there not equal or...? and I think it divides it or something im not fully sure on that lol

OpenStudy (erinweeks):

im trying though lol. almost done this is the last question then I have to get ready for work lol

jimthompson5910 (jim_thompson5910):

we have this